0.58g of flavoring substance is used to improve the taste of the pineapple cake. What is the number of molecules of the flavoring substance? [RMM flavoring substance = 116 gmol-1]


1
Expert's answer
2021-10-13T04:47:43-0400

Solution:

RMM flavoring substance = 116 g mol-1

Hence,

(0.58 g) × (1 mol / 116 g) = 0.005 mol of flavoring substance


One mole of any substance contains 6.02×1023 formula units (Na).

Hence,

(0.005 mol) × (6.02×1023 molecules / 1 mol) = 3.01×1021 molecules of the flavoring substance


Answer: The number of molecules of the flavoring substance is 3.01×1021